# HADIS: Hybrid Adaptive Diffusion Model Serving for Efficient Text-to-Image Generation

### v1

@@ 3. HADIS Design
We present the architecture of HADIS, a diffusion model serving system that incorporates the architectural elements presented in §2. We first provide an overview (§3.1) and then detail the main features of HADIS, including adaptive model cascading (§3.2), hybrid routing (§3.3), and resource management (§3.4).
Figure 4. System Architecture of HADIS.
3.1. Overview of HADIS
@@
Figure 4 shows the system architecture of HADIS. It has three major components: Controller, Query Router, and Workers. We explain how these components are involved in the control path and the data path. The control path is marked with dotted arrows in Figure 4. HADIS is initialized with a set of diffusion models registered in the Controller. The control path takes mainly two actions: (1) The Controller profiles the registered models and pairwise model combinations offline to generate a cascade configuration lookup table. (2) At runtime, the Controller leverages this table to select the optimal cascade configuration. Each configuration includes a model pair and the associated thresholds. The data path is marked with solid arrows in Figure 4. (1) A query from the client is sent to the query router. (2) The router estimates query hardness and routes the query to a worker hosting a suitable diffusion-model variant. (3) If the query is first processed by the lightweight model and the quality score of the generated response is below the threshold, the worker re-routes the query to another worker with a more heavyweight model. (4) The final output is returned to the client. We next describe each main system component.
@@
Controller. A centralized Controller manages resources and coordinates system modules. The Controller includes the following components: (1) A Resource Manager that periodically solves MILP to determine the model pair and thresholds (i.e., the router’s hardness threshold and the discriminator’s quality threshold) and to produce a resource allocation plan (i.e., model placement across workers and batch sizes); (2) A Load Balancer that generates routing tables so each query is directed to a worker, and updates these tables when model selection or placement changes. (3) A Metadata store that periodically pulls heartbeats from the Query Router and Workers to update execution statistics, including per-model queue lengths and demand at each worker. (4) A Model Profiler that supplies a lookup table of profiled cascade configurations to the Resource Manager. Only registered diffusion-model pairs and discriminators are eligible for adaptive model cascading. §3.2 explains construction of the lookup table and §3.4 presents the resource-management formulation based on the lookup table.
@@
3.2. Adaptive Model Cascading
Adaptive model cascading dynamically selects the appropriate model pairs in a cascade given the current demand and latency SLO. However, determining both the optimal cascade and resource-allocation plan is NP-hard. The search space can be prohibitively large due to numerous model candidates, decision thresholds, model placements, and batch-size choices. HADIS addresses this challenge by decomposing the problem into two stages: offline model profiling and online MILP-based resource management.
Model candidate selection. The model pool could contain many variants spanning a broad latency-quality spectrum. Exhaustively profiling all combinations is impractical. We therefore prune the pool with three heuristic, reproducible rules before any offline benchmarking. Figure 2(left) illustrates model selection from a pool. Models colored in orange are retained as candidates, while the blue ones are discarded. The three heuristic rules are:
  • Adjacency pruning. If multiple models exhibit similar latency and FID, we retain a single representative and discard the others.
  • Dominance pruning. If one model Pareto-dominates another (no worse in latency and strictly better in quality, or vice versa), we keep the dominating model and drop the dominated ones.
  • Intermediate redundancy. If a model lies approximately on the segment between two other models that are close in latency, we omit it. In a two-stage cascade, thresholding between the neighbors can reproduce its trade-off.
Offline profiling and lookup-table construction. After selecting candidates, we profile two-model cascades formed by ordered pairs ¡lightweight, heavyweight¿. For each pair we sweep the hardness threshold and quality threshold over a fixed grid. Each sweep point defines a cascade configuration, which we evaluate and record:
@@ 3.2. Adaptive Model Cascading
  • diff_thres: the router’s hardness threshold
  • qual_thres: the discriminator’s quality threshold
  • route_ratio: the measured fraction of queries processed by each model under these thresholds
@@
We then prune dominated rows to retain only Pareto-optimal configurations. The union over all pairs yields a compact lookup table of ¡pair, thresholds¿ configurations that captures the relevant quality–latency trade-offs and induced routing fractions. This table is the sole input to the online phase, where the Resource Manager selects one configuration and jointly decides model placement and batch sizes via a fast MILP, enabling HADIS to adapt model pairs and thresholds at runtime without expensive online exploration.
3.3. Hybrid Routing
HADIS employs: 1) a rule-based router that estimates query hardness before any generation, and 2) a CLIP (Contrastive Language-Image Pre-training)-based discriminator that evaluates the generated image. Hard queries can bypass the lightweight model, while the discriminator catches false “easy” cases after a lightweight model. The decision thresholds are tuned online by the Resource Manager in §3.4.
@@ 3.3. Hybrid Routing
Table 1. Features and proposed rules of the rule-based router.
@@
CLIP-based discriminator. The Discriminator determines whether the output of the lightweight model is aesthetically acceptable. If so, the result is returned to the client, otherwise the query is re-routed to the heavyweight model. We use a CLIP backbone (Radford et al., 2021) because it provides robust image embeddings for aesthetic evaluation (Hentschel et al., 2022). The discriminator keeps the pretrained CLIP frozen and fine-tunes a multi-layer perceptron (MLP) head. The model is trained to distinguish generated images from real images and then used for inference to decide whether a lightweight output meets the quality threshold.
@@
Training. We train the discriminator as a binary classifier to distinguish high-quality, real images (e.g., ImageNet (Russakovsky et al., 2015)) from images generated by our candidate diffusion models. This teaches it to detect typical divergences, such as loss of sharpness, texture incoherence, and structural glitches, enabling accurate quality assessment within the cascade.
@@
Inference. Given a prompt and its generated image, CLIP produces text and visual embeddings; the MLP head consumes the visual embedding and outputs a softmax score in $[0,1]$, used as the quality score (likelihood the image is “real-like” and thus high quality). The discriminator can also compute the CLIP score (Hessel et al., 2021) (cosine similarity between visual and text embeddings) to assess semantic alignment at no extra inference cost. While we do not report semantic metrics in our evaluation because candidates differ little on CLIP score, the signal is useful in production and can be combined with the quality score as a safety check.
Hybrid routing with thresholds. The hardness threshold $\theta$ and quality threshold $\tau$ are the keys to HADIS’s model scaling workflow, as shown in Figure 1. These parameters are co-optimized with other system settings such as model pairs and batch sizes, which are discussed at §3.4. After lightweight inference produces an image, the discriminator outputs a scalar quality score. If the score is below $\tau$, the query is re-routed to the heavyweight stage. Increasing $\tau$ improves response quality but raises compute and latency as more borderline outputs are rejected and re-routed. On the other hand, obviously hard queries can bypass the lightweight stage when the router’s hardness score exceeds $\theta$, saving resources and latency. Increasing $\theta$ makes the router stricter about labeling prompt “hard”, so a larger fraction of queries first pass through the lightweight stage. This may reduce average queueing at the heavyweight stage but risks higher overall system overhead if lightweight outputs subsequently fail the quality criterion determined by $\tau$.
3.4. Resource Management
@@
The Resource Manager adapts HADIS to the current workload by jointly (i) selecting the cascade configuration, including a lightweight/heavyweight model pair and thresholds, (ii) placing models across workers, and (iii) choosing batch size per model. It tunes thresholds (i.e., the router’s hardness threshold and discriminator’s quality threshold) to balance response quality and latency as demand shifts. For instance, during low demand, a lower hardness threshold or higher quality threshold prioritizes image quality, while at peak times, a higher hardness threshold or lower quality threshold ensures fewer queries are routed to the heavyweight models, allowing minor quality compromises to meet latency deadlines. When observed demand exceeds the capacity of the current pair under its active thresholds, the Resource Manager shifts to a lighter configuration and vice versa. These changes are coordinated with model placement and batch-size updates so that workers are reallocated to match the new routing schedule and each model operates near its latency/throughput sweet spot.
@@
Optimization variables. For each model pair, we introduce a binary selector $z_{k}$ over the threshold configuration $k$ of that pair. A threshold configuration $k$ includes a hardness threshold and a quality threshold (i.e., $k=<\tau,\theta>$). For each model $i$ in the pair, we choose an integer worker count $x_{i}\geq 0$ and a one-hot batch size via binaries $y_{i,b}$ over a discrete set of batches $b$. The lookup table provides the routing fraction $r_{k,j}$ to model $i$ under configuration $k$, the profiled latency $L_{i,b}$ and throughput $\mu_{i,b}$ at batch size $b$, and the quality of the configuration FIDk. Let $\lambda$ denote the system demand, $S$ the available workers for this pair, and $T_{SLO}$ the latency deadline. Queueing delay for model $i$ is estimated online from the instantaneous queue length and the arrival rate.
@@
(1) $\sum_{b}y_{i,b}=1,\quad\sum_{k}z_{k}=1,\quad\sum_{i}x_{i}\leq S\quad\forall i$
@@
(2) $x_{i}\sum_{b}y_{i,b}\mu_{i,b}\geq\lambda\sum_{k}z_{k}r_{k,i}\quad\quad\forall i$
Then, considering the latency constraint, end-to-end latency must remain within the SLO. Let the selected batch latency for model $i$ be $\widehat{L}_{i}=\sum_{b}y_{i,b}L_{i,b}$ and let $D^{\text{queue}}_{i}=\alpha\frac{Q_{i}}{\lambda_{i}}$ denote the queueing delay estimate given Little’s law (Shortle et al., 2018) where $\alpha$ is a safety factor. We impose a conservative path bound, making sure the latency constraint is met:
(3) $\sum_{i}(\widehat{L}_{i}+D^{\text{queue}}_{i})\leq T_{SLO}$
@@ 4. Evaluation
4.1. Experiment settings
Implementation of HADIS. We implement HADIS on a testbed cluster to evaluate performance on real GPU hardware. The system implementation consists of $\sim$5K lines of Python code. We use the HuggingFace (von Platen et al., 2022) and PyTorch frameworks (Paszke et al., 2019) to execute the diffusion models and discriminator, and use spaCy (Honnibal et al., 2020) to extract prompt features for the rule-based query router. Our cluster consists of 16 workers with NVIDIA L40S GPU, and system components including Controller, Load Balancer, Worker, and Client, communicate via gRPC for low-latency coordination. We use Gurobi (Gurobi, 2024) to solve our MILP optimization. Unless otherwise noted, all results reported in this paper are collected on this testbed.
Model Selection. We apply our model-selection rules to choose four diffusion models that provide diverse latency-quality trade-offs, including SDXL-Lightning, SD3.5-Turbo, SD3.5-Medium, and SD3.5-Large. SDXL-Lightning (Lin et al., 2024) is a fast generative model that produces an image in a few steps; in our setup, we use 2 steps, yielding $\sim$0.5s latency per $\times1024$ image on an L40S GPU. Stable Diffusion 3.5 (AI, 2024) is a diffusion transformer family that can produce images of high quality. We configure SD3.5-Turbo with 4 steps ($\sim$1.3s per image on L40S), while SD3.5-Medium and SD3.5-Large use 50 steps, with latencies of $\sim$13s and $\sim$27s, respectively, on the same hardware. Unless otherwise noted, the latency SLO is 60s, chosen as a multiple of the slowest model’s runtime, and all models render at $\times1024$ resolution.
Datasets and workloads. For quality evaluation, we use ImageNet-1k, a widely-used subset of ImageNet (Russakovsky et al., 2015), with associated captions. We take the first 5k caption-image pairs, using captions as prompts and the corresponding images as the reference set for FID computation. To drive system load, we use the Microsoft Azure Functions trace (Shahrad et al., 2020) as a representative real-world workload, applying shape-preserving scaling to match our cluster’s capacity.
@@ 4.1. Experiment settings
Baselines. We compare HADIS against a spectrum of baselines ranging from fully static (Clipper) to dynamic resource allocation (Proteus) and query-aware cascading (DiffServe).
  • Clipper (Crankshaw et al., 2017a) is a static baseline, which does not dynamically select models or allocate resources, and cannot route queries given query complexity. We implement two versions of Clipper, Clipper-Light, minimizing SLO violations with the most lightweight model (SDXL-Lightning), and Clipper-Heavy, maximizing response quality with the most heavyweight model (SD3.5-Large). Note that Clipper is also representative of other static inference-serving systems, such as TensorFlow-Serving (Olston et al., 2017), which depend on the application developer to handle the resource allocation of models.
  • Proteus (Ahmad et al., 2024) is a high-throughput serving system that selects among multiple model variants as workload changes. It scales quality by switching to faster variants from the four diffusion model candidates under high load and switching to higher-quality variants when capacity allows. Proteus runs single-stage inference without a per-query cascade or discriminator, where queries are randomly routed to each worker with equal probability.
  • DiffServe (Ahmad et al., 2025) is a text-to-image serving system built around a model-cascading pipeline. DiffServe uses an MILP planner to adapt to demand by jointly setting resource allocations and the quality threshold of the discriminator. Unlike HADIS, it employs a fixed pair of light and heavyweight models. In our implementation, we fix the pair to SD3.5-Turbo and SD3.5-Large, which outperformed other pairings on the query workload.
@@
Figure 6. Performance of HADIS on synthetic traces. (a) Time-series of a piecewise-constant demand trace and system performance. Top: demand QPS with four phases (P1-P4). Middle: FID rises when demand increases and recovers as demand decreases. Bottom: SLO violation ratio stays low overall, with brief decreases during P3/4 and increases during P1/2 due to the latency of models in use. (b) Left: queries are shifted toward faster variants under higher load and back toward higher-quality variants as load decreases. Right: number of workers assigned to lightweight/heavyweight models is coordinated with the ratio of queries routed to heavyweight models to maintain low SLO violations while maximizing quality as demand fluctuates.
4.2. Performance on Real-World Traces

### v2

@@ 4 HADIS Design
We present HADIS, a diffusion model serving system that incorporates the architectural elements presented in §2. We first provide an overview (§4.1), then detail the main features of HADIS, including adaptive model cascading (§4.2), hybrid routing (§4.3), and resource management (§4.4).
4.1 Overview of HADIS
Figure 6: System Architecture of HADIS.
@@ 4.1 Overview of HADIS
Fig. 6 shows the system architecture of HADIS. It has three major components: Controller, Query Router, and Workers. We explain how these components are involved in the control path and the data path. The control path is marked with dotted arrows in Fig. 6. HADIS is initialized with a set of diffusion models registered in the Controller. The control path takes mainly two actions: (1) The Controller profiles the registered models and pairwise model combinations offline to generate a cascade configuration lookup table. (2) At runtime, the Controller leverages this table to select the optimal cascade configuration. Each configuration includes a model pair and the associated thresholds. The data path is marked with solid arrows in Fig. 6. (1) A query from the client is sent to the query router. (2) The router estimates query hardness and routes the query to a worker hosting a suitable diffusion-model variant. (3) If the query is first processed by the lightweight model and the quality score of the generated response is below the threshold, the worker re-routes the query to another worker with a more heavyweight model. (4) The final output is returned to the client. We next describe each main system component.
@@
Controller. A centralized Controller manages resources and coordinates system modules. It comprises: (1) A Resource Manager that periodically solves MILP to determine the model pair and thresholds (i.e., the router’s hardness threshold and the discriminator’s quality threshold) and to produce a resource allocation plan (i.e., model placement across workers and batch sizes); (2) A Load Balancer that generates routing tables so each query is directed to a worker, and updates these tables when model selection or placement changes. (3) A Metadata store that periodically pulls heartbeats from the Query Router and Workers to update execution statistics, including per-model queue lengths and demand at each worker. (4) A Model Profiler that supplies a lookup table of profiled cascade configurations to the Resource Manager. Only registered diffusion-model pairs and discriminators are eligible for adaptive model cascading. §4.2 describes construction of the lookup table and §4.4 presents the resource-management formulation based on the lookup table.
@@
4.2 Adaptive Model Cascading
Adaptive model cascading dynamically selects the appropriate model pairs in a cascade given the current demand and latency SLO. However, determining both the optimal cascade and resource-allocation plan is NP-hard. The search space can be prohibitively large due to numerous model candidates, decision thresholds, model placements, and batch-size choices. HADIS addresses this challenge by decomposing the problem into two stages: offline model profiling (this section) and online MILP-based resource management (\$4.4).
Model candidate selection. In practice, the model pool is already constrained by user provision, so we do not face an unbounded number of variants. However, exhaustively profiling every possible cascade built from these models is still unnecessary. We therefore prune the pool with three heuristic, reproducible rules before any offline benchmarking: (1) If multiple models exhibit similar latency and FID, we retain a single representative and discard the others; (2) If one model Pareto-dominates another (no worse in latency and strictly better in quality, or vice versa), we keep the dominating model and drop the dominated ones; (3) If a model lies approximately on the line segment between two neighbors, we omit it, as a two-stage cascade can reproduce its trade-off by thresholding between the neighbors. Figure 2-left illustrates the retained (orange) and pruned (blue) models. In our running example, we start with 8 model candidates, yielding 28 two-model cascades and 2800 configurations (10 hardness thresholds and 10 quality thresholds per cascade). After pruning 4 models, only 4 remains, yielding 6 cascades and 600 configurations, reducing profiling cost by 79%.
Offline profiling and lookup-table construction. The theoretical analysis in §3 characterizes which router–discriminator operating points would be optimal if we knew the true easy/hard split and exact ROC curves. However, in practice there is no ground-truth notion of “easy vs. hard” or per-query correctness label for text-to-image tasks, so we cannot directly measure $q$, TPR, and FPR as in the analysis. Instead, after determining the two-model cascade <lightweight, heavyweight>, for each pair we sweep the hardness threshold and quality threshold over a fixed grid to form a cascade configuration:
@@ 4.2 Adaptive Model Cascading
  • diff_thres: the router’s hardness threshold ($\theta$)
  • qual_thres: the discriminator’s quality threshold ($\tau$)
  • route_ratio: the measured fraction of queries processed by each model under these thresholds ($r_{0}$ and $r_{1}$)
@@
We then prune dominated rows to retain only Pareto-optimal configurations. The union over all pairs yields a compact lookup table of <pair, thresholds> configurations that captures the relevant quality–latency trade-offs and induced routing fractions. This table is the sole input to the online phase, where the Resource Manager jointly select a configuration, model placement, and batch sizes via a fast MILP, enabling HADIS to adapt model pairs and thresholds at runtime without expensive online exploration.
4.3 Hybrid Routing
HADIS employs: (1) a rule-based router that estimates query hardness before any generation, and (2) a CLIP (Contrastive Language-Image Pre-training)-based discriminator that evaluates the generated image. While an explicit ROC to characterize the performance of the router and discriminator is infeasible due to the lack of ground truth labels, we empirically show that the design enables HADIS to outperform non-hybrid alternatives in §5.4.
@@ 4.3 Hybrid Routing
Table 1: Features and proposed rules of the rule-based router.
@@
CLIP-based discriminator. The Discriminator is co-located with lightweight models on GPUs and determines whether their outputs are aesthetically acceptable. Accepted results are returned to the client, while rejected ones are re-routed to heavyweight models. We use a CLIP backbone [33] because it provides robust image embeddings for aesthetic evaluation [20]. The discriminator keeps the pretrained CLIP frozen and fine-tunes a multi-layer perceptron (MLP) head.
@@
Training. We train the discriminator as a binary classifier to distinguish high-quality, real images (e.g., ImageNet [35]) from images generated by our candidate diffusion models. This teaches it to detect typical divergences, such as loss of sharpness, texture incoherence, and structural glitches, enabling accurate quality assessment within the cascade.
@@
Inference. Given a prompt and its generated image, CLIP produces text and visual embeddings; the MLP head consumes the visual embedding and outputs a softmax score in $[0,1]$, used as the quality score (likelihood the image is “real-like” and thus high quality). The discriminator can also compute the CLIP score [21] (cosine similarity between visual and text embeddings) to assess semantic alignment at no extra inference cost. While we do not report semantic metrics in our evaluation because candidates differ little on CLIP score, the signal is useful in production and can be combined with the quality score as a safety check. The overhead of discriminator is analyzed at §5.8.
4.4 Resource Management
@@
The Resource Manager adapts HADIS to the current workload by jointly (i) selecting the cascade configuration, including a lightweight/heavyweight model pair and thresholds, (ii) placing models across workers, and (iii) choosing batch size per model. It tunes thresholds (i.e., the router’s hardness threshold and discriminator’s quality threshold) to balance response quality and latency as demand shifts. For instance, during low demand, a lower hardness threshold or higher quality threshold prioritizes image quality, while at peak times, a higher hardness threshold or lower quality threshold ensures fewer queries are routed to heavyweight models, allowing minor quality compromises to meet latency deadlines. When demand exceeds the capacity of the current pair under its active thresholds, the Resource Manager shifts to a lighter configuration and vice versa. These changes are coordinated with model placement and batch-size updates so that workers are reallocated to match the new routing schedule and each model operates near its latency/throughput sweet spot.
@@
Optimization variables. For each model pair, we introduce a binary selector $z_{k}$ over the threshold configuration $k$ of that pair. A threshold configuration $k$ includes a hardness threshold and a quality threshold (i.e., $k=<\tau,\theta>$). For each model $i$ in the pair, we choose an integer worker count $x_{i}\geq 0$ and a one-hot batch size via binaries $y_{i,b}$ over a discrete set of batches $b$. The lookup table provides the routing fraction $r_{k,i}$ to model $i$ under configuration $k$, the profiled latency $L_{i,b}$ and throughput $\mu_{i,b}$ at batch size $b$, and the quality of the configuration FIDk. Let $\lambda$ denote the system demand, $S$ the available workers for this pair, and $T_{SLO}$ the latency deadline. Queueing delay for model $i$ is estimated online from the instantaneous queue length $Q_{i}$ and the arrival rate $\lambda_{i}$.
@@
$\sum_{b}y_{i,b}=1,\quad\sum_{k}z_{k}=1,\quad\sum_{i}x_{i}\leq S\quad\forall i$ (3)
@@
$x_{i}\sum_{b}y_{i,b}\mu_{i,b}\geq\lambda\sum_{k}z_{k}r_{k,i}\quad\quad\forall i$ (4)
Then, considering the latency constraint, end-to-end latency must remain within the SLO. Let the selected batch latency for model $i$ be $\widehat{L}_{i}=\sum_{b}y_{i,b}L_{i,b}$ and let $D^{\text{queue}}_{i}=\alpha\frac{Q_{i}}{\lambda_{i}}$ denote the queueing delay estimate given Little’s law [40] where $\alpha$ is a safety factor. We impose a conservative path bound, making sure the latency constraint is met:
$\sum_{i}(\widehat{L}_{i}+D^{\text{queue}}_{i})\leq T_{SLO}$ (5)
@@ 5 Evaluation
5.1 Experiment settings
Implementation of HADIS. We implement HADIS on a testbed cluster to evaluate performance on real GPU hardware. The system implementation consists of $\sim$5K lines of Python code. We use the HuggingFace [41] and PyTorch frameworks [30] to execute the diffusion models and discriminator, and use spaCy [23] to extract prompt features for the rule-based query router. Our cluster consists of 16 workers with NVIDIA L40S GPU, and system components including Controller, Load Balancer, Worker, and Client, communicate via gRPC for low-latency coordination. We use Gurobi [19] to solve our MILP optimization. Unless otherwise noted, all results reported in this paper are collected on this testbed.
Model Selection. We apply our model-selection rules to choose four diffusion models that provide diverse latency-quality trade-offs, including SDXL-Lightning, SD3.5-Turbo, SD3.5-Medium, and SD3.5-Large. We use the default number of steps for each model, ensuring high-quality generation. SDXL-Lightning [26] is a fast generative model that produces an image in 2 steps, yielding $\sim$0.5s latency per $\times1024$ image on an L40S GPU. Stable Diffusion 3.5 [9] is a diffusion transformer family that can produce images of high quality. We configure SD3.5-Turbo with 4 steps ($\sim$1.3s per image on L40S), while SD3.5-Medium and SD3.5-Large use 50 steps, with latencies of $\sim$13s and $\sim$27s, respectively, on the same hardware. Unless otherwise noted, the latency SLO is 60s, chosen as a multiple of the slowest model’s runtime, and all models render at $\times1024$ resolution.
Datasets and workloads. For quality evaluation, we use ImageNet-1k, a widely-used subset of ImageNet [35], with associated captions. We take the first 5k caption-image pairs, using captions as prompts and the corresponding images as the reference set for FID computation. To drive system load, we use the Microsoft Azure Functions trace [38] as a representative real-world workload, applying shape-preserving scaling to match our cluster’s capacity.
@@ 5.1 Experiment settings
Baselines. We compare HADIS against a spectrum of baselines ranging from static (Clipper) to dynamic resource allocation (Proteus) and query-aware cascading (DiffServe).
  • Clipper [13] is a static baseline, which does not dynamically select models, allocate resources, or route queries given query complexity. We implement two variants, Clipper-Light, which minimizes SLO violations with the lightest model (SDXL-Lightning), and Clipper-Heavy, which maximizes response quality with the heaviest model (SD3.5-Large). Note that Clipper also represents other static serving systems (e.g., TensorFlow-Serving [28]), which rely on application developers to allocate resources for models.
  • Proteus [7] is a high-throughput serving system that selects among multiple model variants as workload changes. It scales quality by switching to faster variants from the four diffusion model candidates under high load and switching to higher-quality variants when capacity allows. Proteus runs single-stage inference without a per-query cascade or discriminator, where queries are randomly routed to each worker with equal probability.
  • DiffServe [8] is a text-to-image serving system built around a model-cascading pipeline. DiffServe uses an MILP planner to adapt to demand by jointly setting resource allocations and the quality threshold of the discriminator. Unlike HADIS, it employs a fixed pair of light and heavyweight models. In our implementation, we fix the pair to SD3.5-Turbo and SD3.5-Large, which outperformed other pairings on the query workload.
@@
Figure 8: Performance of HADIS on synthetic traces. (a) Time-series of a piecewise-constant demand trace and system performance. Top: demand QPS with four phases (P1-P4). Middle: FID rises when demand increases and recovers as demand decreases. Bottom: SLO violation ratio stays low overall, with brief decreases during P3/4 and increases during P1/2 due to the latency of models in use. (b) Left: queries are shifted toward faster variants under higher load and back toward higher-quality variants as load decreases. Right: number of workers assigned to lightweight/heavyweight models is coordinated with the ratio of queries routed to heavyweight models to maintain low SLO violations while maximizing quality as demand fluctuates.
5.2 Performance on Real-World Traces